You have a subject with the following pulmonary function study results. TLC = 7 L, VC = 2.2 L, IC = 4.7 L, VT = 0.5L. Calculate missing lung volumes or capacities and answer related questions.

Find ERV, FRC, IC
Answer #1

Residual volume (RV) = Total lung capacity (TLC) - Vital capacity (VC)

                                    = 7 - 2.2

                              RV = 4.8 L

Expiratory Reserve Volume (ERV) = Inspiratory capacity (IC) - Residual volume (RV)

                                                       = 4.7 - 4.8

                                              ERV = 0.1 L

Functional residual capacity (FRC) = ERV + RV

                                                        = 0.1 + 4.8

                                                FRC = 4.9 L

Inspiratory reserve volume (IRV) = Vital capacity (VC) - Tidal volume (TV) - Expiratory Reserve Volume (ERV)

                                                     = 2.2 - 0.5 - 0.1

                                              IRV = 1.6 L

Inspiratory capacity (IC) = TV + IRV

                                       = 0.5 + 1.6

                                  IC = 2.1 L